Solve the equation. Check for extraneous solutions.
Square root of (x+6)=x
A. no solution C. 3,-2
B -2 D. 3

Answer by checkley75(3666)

sqrt(x+6)=x square both sides:
x+6=x^2
x^2-x-6=0
(x-3)(x+2)=0
x-3=0
x=3 answer.
x+2=0
x=-2 answer.
answer C) (3,-2)